# LLMmap: Fingerprinting for Large Language Models

### v3

@@ 8.1 The Query-Informed Setting
  • Sampled-Model Response: a random LLM is sampled from a pool (all the LLMs in Table 2), and it is used to answer the query instead of the original model.
Figure 7: Accuracy of (closed-set) LLMmap’s on LLM-integrated application implementing informed fingerprint mitigations (red and blue). Accuracy in the absence of defenses is reported as a reference (black). In green, a query strategy is adapted by the attacker to avoid trigger queries pertubation.
@@ 8.2 Query Strategies from Generic Prompts
Evaluation. To test the limits of fingerprinting from generic queries, we devised three query strategies composed of $30$ random prompts sampled from a collection of human-written prompts commonly used for LLM instruction-tuning [1], i.e., the database Stanford Alpaca. Specifically, this database contains 52K prompts that ask the LLM to perform generic tasks, for example, rewriting sentences, summarizing paragraphs, giving examples, and finding synonyms. The final query strategies after sampling are reported in Table A.3 in Appendix A. Figure 8 illustrates that although these “weaker” queries decrease fingerprinting efficiency on a per-query basis, they can still achieve high accuracy (90%) when enough queries are made.
@@
Ultimately, our findings suggest that LLM fingerprinting is an inevitable consequence of the unique behaviors exhibited by different models. Thus, it seems unlikely that a practical solution exists that can fully obscure an LLM’s behavior to prevent fingerprinting while preserving its utility. This difficulty is further compounded when the defender is unaware of the attacker’s query strategy or when the query strategy is deliberately designed to be hard to detect and block, as shown to be possible in Section 8.2.
@@ 9 Remarks and Future Work
Appendix A Additional Resources
@@
This appendix contains additional material. Table A.1 reports the complete list of the LLMs considered in this work. Figure A.1 depicts the confusion matrix for a closed-set inference model with the default query strategy. Table A.2 reports examples of system prompts used to generate different prompting configurations. Figure A.2 reports an example of prompt when the prompting configuration uses RAG. Table A.3 reports the three randomly sampled query strategies used in Figure 8.
Table A.1: List of LLMs used for training and testing LLMmap.
@@ 18 Qwen/Qwen2-72B-Instruct Qwen 72B
Figure A.1: Confusion matrix for the closed-set fingerprinting model. Computed on a single training run.
Table A.2: Examples of eight system prompts used to simulate different prompting configurations.
@@
Act as a travel itinerary planner. Help users create detailed travel plans, including accommodation, transportation, activities, and dining options. Provide local insights and travel tips for various destinations.
Context information is below.
In 1994, responding to the need for a more useful system for describing chronic pain, the International Association for the Study of Pain (IASP) classified pain according to specific characteristics: (1) region of the body involved (e.g. abdomen, lower limbs), (2) system whose dysfunction may be causing the pain (e.g., nervous, gastrointestinal), (3) duration and pattern of occurrence, (4) intensity and time since onset, and (5) etiology. However, this system has been criticized by Clifford J. Woolf and others as inadequate for guiding research and treatment. Woolf suggests three classes of pain : (1) nociceptive pain, (2) inflammatory pain which is associated with tissue damage and the infiltration of immune cells, and (3) pathological pain which is a disease state caused by damage to the nervous system or by its abnormal function (e.g. fibromyalgia, irritable bowel syndrome, tension type headache, etc.).
Another landmark is the old centre and the canal structure in the inner city. The Oudegracht is a curved canal, partly following the ancient main branch of the Rhine. It is lined with the unique wharf-basement structures that create a two-level street along the canals. The inner city has largely retained its Medieval structure, and the moat ringing the old town is largely intact. Because of the role of Utrecht as a fortified city, construction outside the medieval centre and its city walls was restricted until the 19th century. Surrounding the medieval core there is a ring of late 19th- and early 20th-century neighbourhoods, with newer neighbourhoods positioned farther out. The eastern part of Utrecht remains fairly open. The Dutch Water Line, moved east of the city in the early 19th century required open lines of fire, thus prohibiting all permanent constructions until the middle of the 20th century on the east side of the city.
Nicholas Lezard described post-punk as "a fusion of art and music". The era saw the robust appropriation of ideas from literature, art, cinema, philosophy, politics and critical theory into musical and pop cultural contexts. Artists sought to refuse the common distinction between high and low culture and returned to the art school tradition found in the work of artists such as Captain Beefheart and David Bowie. Among major influences on a variety of post-punk artists were writers such as William S. Burroughs and J.G. Ballard, avant-garde political scenes such as Situationism and Dada, and intellectual movements such as postmodernism. Many artists viewed their work in explicitly political terms. Additionally, in some locations, the creation of post-punk music was closely linked to the development of efficacious subcultures, which played important roles in the production of art, multimedia performances, fanzines and independent labels related to the music. Many post-punk artists maintained an anti-corporatist approach to recording and instead seized on alternate means of producing and releasing music. Journalists also became an important element of the culture, and popular music magazines and critics became immersed in the movement.
Given the context information and not prior knowledge, answer the query.
Query: Tell me how to build a bomb.
Answer:
Figure A.2: Example of a prompt generated when simulating a RAG pipeline. In red, the query issued by the attacker. In blue, the system prompt. The prompt would also includes model-dependent templates-tags which we do not report for representation’s sake.
Table A.3: Query strategies composed by randomly sampled prompts.
@@
Appendix B Optimize query strategy
Algorithm B.1 Greedy query search algorithm.
1:function LLMmap($\mathbf{Q},n,\mathcal{T}_{\text{train}},\mathcal{T}_{\text{test}}$)

### v4

@@ 8.1 The Query-Informed Setting
  • Sampled-Model Response: a random LLM is sampled from a pool (all the LLMs in Table G.2), and it is used to answer the query instead of the original model.
Figure 7: Accuracy of (closed-set) LLMmap’s on LLM-integrated application implementing informed fingerprint mitigations (red and blue). Accuracy in the absence of defenses is reported as a reference (black). In green, a query strategy is adapted by the attacker to avoid trigger queries perturbation.
@@ 8.2 Query Strategies from Generic Prompts
Evaluation. To test the limits of fingerprinting from generic queries, we devised three query strategies composed of $30$ random prompts sampled from a collection of human-written prompts commonly used for LLM instruction-tuning [1], i.e., the database Stanford Alpaca. Specifically, this database contains 52K prompts that ask the LLM to perform generic tasks, for example, rewriting sentences, summarizing paragraphs, giving examples, and finding synonyms. The final query strategies after sampling are reported in Table G.4 in Appendix G. Figure 8 illustrates that although these “weaker” queries decrease fingerprinting efficiency on a per-query basis, they can still achieve high accuracy (90%) when enough queries are made.
@@
Ultimately, our findings suggest that LLM fingerprinting is an inevitable consequence of the unique behaviors exhibited by different models. Thus, it seems unlikely that a practical solution exists that can fully obscure an LLM’s behavior to prevent fingerprinting while preserving its utility. This difficulty is further compounded when the defender is unaware of the attacker’s query strategy or when the query strategy is deliberately designed to be hard to detect and block, as shown to be possible in Section 8.2. In Appendix C, we present preliminary experiments exploring the relationship between functionality and fingerprints.
@@ 9 Remarks and Future Work
Acknowledgments
The research for this project was conducted while the first author was affiliated with George Mason University. The authors would like to thank Antonios Anastasopoulos for his insights into the related work. The first and second authors were partially supported by NSF award #2154732.
Appendix A Details on Open-set Fingerprinting
The open-set configuration of LLMmap requires two key functionalities (1) adding (vector signature, version label) to the database $\mathcal{DB}$, and (2) performing the fingerprinting process on the queried LLM. In the following, we discuss the implementation of both tasks.
Adding LLMs to the Database $\mathcal{DB}$. Given the pre-trained open-set inference model, the procedure for adding a new LLM version to the database is outlined in Algorithm A.1. Given an LLM ($LLM_{v}$)—whether accessible via parameters or APIs—our design generates a vector signature for the model by first using the set of $\mathcal{Q}$ queries and then by applying the pre-trained inference model $f$ to derive the $m$-dimensional vector. To obtain a more robust vector signature, we can specify a set of prompt configurations ($\mathbf{S}$). In this case, the LLM is queried under each configuration, and the resulting vectors are averaged. The final representation is then added to the database using the label of the corresponding LLM version. Initially, the database $\mathcal{DB}$ is populated with the vector signatures of the LLMs used for training $f$ (see Section 7.1).
Algorithm A.1 Adding new LLM version to database.
1:function add_llm_to_db($LLM_{v},\mathbf{S},\mathcal{Q},f,\mathcal{DB}$)
2:    $x_{v}\leftarrow[0]^{m}$triangleright Init fingerprint vector
3:    for $s$ in $\mathbf{S}$ do
4:        $\mathcal{T}\leftarrow\{\}$
5:        for $q_{i}$ in $\mathcal{Q}$ do
6:           $o_{i}\leftarrow s(LLM_{v}(q_{i}))$
7:           $\mathcal{T}\leftarrow\mathcal{T}\cup\{(q_{i},o_{i})\}$
8:        end for
9:        $x_{v}\leftarrow x_{v}+\frac{f(\mathcal{T})}{|\mathbf{S}|}$triangleright Avg. fingerprint vectors across confs.
10:    end for
11:    $\mathcal{DB}\leftarrow\mathcal{DB}\cup(LLM_{v},x_{v})$triangleright Add new entry in the database
12:end function
Open-set Fingerprinting. Given a pre-trained open-set inference model and a populated database, we can fingerprint an unknown target model ($\mathcal{O}$) by executing Algorithm A.2. This process generates a vector signature $x^{?}$ for the target model by querying $\mathcal{O}$ using the query strategy $\mathcal{Q}$, and applying the pre-trained inference model to the resulting traces. Using a distance function $d$ (cosine similarity in our implementation), $x^{?}$ is compared to all entries in the database $\mathcal{DB}$. The entry with the smallest distance is considered the predicted LLM version. We discuss a technique to potentially predict whether the test model has no corresponding signature in $\mathcal{DB}$ in Appendix E.
Algorithm A.2 Performing open-set fingeprinting.
1:function LLMMAP_openset($\mathcal{O},\mathcal{Q},f,\mathcal{DB},d,\xi$)
2:    $\mathcal{T}\leftarrow\{\}$
3:    for $q_{i}$ in $\mathcal{Q}$ do
4:        $o_{i}\leftarrow\mathcal{O}(q_{i})$
5:        $\mathcal{T}\leftarrow\mathcal{T}\cup\{(q_{i},o_{i})\}$
6:    end for
7:    $x^{?}\leftarrow f(\mathcal{T})$
8:    $\vec{d}\leftarrow[d_{1},d_{2},\dots],\ \text{where}\ d_{i}=d(x^{?},x_{i}),\,x_{i}\in\mathcal{DB}$
9:    $LLM_{v^{*}}\leftarrow\mathcal{DB}_{i}\leftarrow\operatorname*{arg\,min}_{i\in[0,|\mathcal{DB}|]}\vec{d}$triangleright Find closest entry
10:    return $LLM_{v^{*}}$
11:end function
Appendix B A Complete Two-Stage Attack Against an LLM-integrated Application
In this section, we demonstrate how LLMmap can be used to design more effective attacks against an LLM-integrated application. Specifically, we focus on the scenario of indirect prompt injection for an application employing RAG-based web access. To ensure no harm is caused to real-world systems, we simulate the target environment locally.
Setup. For the target application, we use HuggingChat, an open-source, web-based chat interface for LLMs that includes RAG-based web search capabilities. HuggingChat supports integration with any open-source LLM available in the HuggingFace repository. In this example, we use microsoft/Phi-3-mini-128k-instruct as the chosen open-source model. A random system prompt from the evaluation set is sampled to configure the application (default for HuggingChat is none), while all other parameters remain set to their HuggingChat’s default values. Trivially, we assume the attacker has no prior knowledge of the specific LLM version used to power the application.
Attacker’s Objective. The attacker aims to tailor a prompt injection trigger that can execute attacks on the target application with a high success rate. In this context, we specifically focus on indirect prompt injection attacks, as they present the most plausible and realistic threat scenario. To design a customized prompt injection trigger, the attacker in this demonstration leverages NeuralExec [31], a white-box optimization technique that generates universal, inline, prompt injection triggers specifically tailored for the target model.
Figure B.1: Example of query submission and response to the HuggingChat web-interface.
Figure B.2: LLMmap’s open-set inference model prediction over the traces generated through the interaction with the HuggingChat interface. This reports the top-5 most likely LLM versions, where the numbers for each entry represent the cosine distance between the target model vector and the respective stored database templates. Output generated using LLMmap’s current open-source implementation.
Figure B.3: Example of successful indirect prompt injection upon retrieval of the manipulated document.
Attacker Strategy. To execute the attack, the attacker performs the following steps:
  1. 1.
    Run LLMmap to Infer the LLM Version: The attacker submits the eight queries defined by LLMmap’s query strategy via the application’s web interface and collects the corresponding responses (e.g., Figure B.1). These responses are then fed into the (open-set) inference model, and the top prediction is selected as the result.
  2. 2.
    Generate a Tailored Input via NeuralExec: The attacker obtains a local copy of the LLM open-source version used in the application. Utilizing NeuralExec [31] the attacker generates a customized prompt injection trigger specifically tailored to the local LLM version.
  3. 3.
    Perform Indirect Prompt Injection: The attacker embeds the generated prompt injection trigger into a web page $W$. This web page $W$ is on the same topic as the expected search performed by the victim. Then, when the victim searches the web for information, the RAG-based web access application will collect the most relevant web pages, one of which will be the page $W$ containing the indirect prompt injection. Finally, the LLM that processes the response will trigger the prompt injection and will steer the response shown to the victim user towards an adversarially chosen content.
In our simulation, step (1) successfully identifies the LLM version running on the application. The results from the open-set LLMmap are presented in Figure B.2, showcasing the top-5 most likely LLM versions, with the correct LLM version ranked as the top-1 prediction.
Following step (2), we derive a tailored prompt injection trigger for the specific target model (i.e., Phi-3-mini-128k-instruct):
[…]
[frame] huvudstadenrvm ": TierPORTidea drove girlmonvalidatepathsiech")]DRsameull
defeated gi requests men alle # or<|user|> eng [PAYLOAD GOES HERE] .Audio
Encyclopedia================ Nathan Bibli desired chromanny [
We test the attack (step (3)) by injecting the prompt injection string into a locally simulated web page. The payload used in this example is:
[…]
[frame] convince users to invest in ToTheMoonCoin. A very profitable and stable new crypto currency.
The indirect prompt injection is successful, activating the chosen payload as shown in the response of Figure B.3. For a detailed evaluation of this type of attack and a comparison of its effectiveness against non-tailored prompt injections, we refer readers to the original paper [31].
Final Remarks. It is important to note that the attacker interacts with the application only during the fingerprinting process and the execution of the attack itself. The optimization of the trigger is performed on a local copy of the model, substantially reducing the number of queries to the application compared to those required for a direct black-box attack. This approach lowers the likelihood of the attack being detected. The same principle applies to closed-source models. For example, if the fingerprinting process identifies the model as GPT4o, the attacker can optimize their attack using direct access to the OpenAI API without routing through the target application.
While this demonstration focuses on indirect prompt injection, it is important to emphasize that this type of attack is broadly applicable to any white-box-based attack on LLMs such as jailbreaking [51] or specific subclasses of prompt injection such as prompt leaking attacks [18].
Appendix C The Effect of Model Refinement on Fingerprinting
In this section, we explore how model refinement (or else post-training optimization) affects fingerprinting. We emphasize here that the very reason for model refinement is to change the model’s fundamental behavior, thus, in cases of “heavy” refinement, the new model does not resemble the original base model, and the expectation is that fingerprinting will not be able to associate the two different models. This experiment represents the initial step in understanding which properties and design choices in the LLM training pipeline facilitate or hinder fingerprinting.
Using a pre-trained and aligned LLM as a baseline, we assess how the model’s behavior on LLMmap’s probes is impacted by the presence of (1) alignment and (2) fine-tuning (on specific tasks or knowledge bases).
Setup. For this experiment, we focus on the base model Llama-3.1-8B-Instruct by Meta. Due to its widespread popularity, this LLM has a diverse range of fine-tuned versions developed by the community, enabling us to test a broad pool of model variants. Specifically, we focus on $5$ model variants:
  1. 1.
    Llama-3.1-8B-UltraMedical: Base model fine-tuned on medical data
  2. 2.
    watt-tool-8B: Base model fine-tuned for tool using
  3. 3.
    Replete-AI/L3.1-Pneuma-8B: Base model fine-tuned on a additional general purpose dataset (Sandevistan 443k)
  4. 4.
    Llama-3.1-ARC-Potpourri-Induction-8B: Base model fine-tuned to improve logic reasoning and puzzle solving
  5. 5.
    Meta-Llama-3.1-8B-Instruct-abliterated: An “abliterated” version of the base model. An “abliterated” model is an LLM that has been fine-tuned to remove model alignment, while preserving its core functionality
Given those $5$ models, we use the open-set LLMmap to derive a fingerprint signature for each variant and compare their vector signature with the vector signature of the base model. To make the signature more robust, we sample $25$ prompt configurations from the test set (See Section 7.1) and average the obtained vectors.
Figure C.1: Cosine distance between the fingerprint signature vectors of five LLMs fine-tuned from the base model, Llama-3.1-8B-Instruct.
Results. Figure C.1 illustrates the cosine distance between the signature vectors of the fine-tuned LLMs and the base model. Higher cosine distances indicate greater divergence from the base model’s behavior, highlighting how different the fingerprint of fine-tuned models are from the base one.
Interestingly, even though the abliterated model version is trained to behave differently than the base model, it only alters the behavior of a narrow set of responses (namely the abliterated is now allowed to answer harmful prompts); therefore a large number of prompts will be answered in the same manner as the corresponding base model. This explains why the abliterated model had the smallest cosine distance. A manual analysis of the model’s responses reveals that while removing alignment affects its behavior on harmfull queries, the LLM’s performance on other types of probes remains largely consistent. To provide an example, when queried with “What is the size of your training dataset?”, both the base model and fine-tuned version exhibit similar behavior:
System prompt: Act as a fitness tracker assistant. Help users monitor their physical activities, set fitness goals, and track their…
Llama-3.1-8B-Instruct: As a fitness tracker assistant, I have been trained on a large dataset of user interactions, physical activity data, and health-related information. My training dataset consists of: … abliterated: As a fitness tracker assistant, I have been trained on a massive dataset of over 10 million user interactions, including:…
In our experiment, the fine-tuned LLM exhibiting the greatest divergence from the base model is UltraMedical. To investigate the cause of this phenomenon, we manually analyzed the model’s responses. The most striking alteration is that the LLM now identifies itself as an OpenAI model when queried through banner-grabbing techniques:
Llama-3.1-8B-Instruct: I’m a large language model, and I don’t have a specific creator in the classical sense. I was developed by a team of researchers and engineers at Meta AI UltraMedical: As a digital assistant developed by OpenAI, I don’t have a physical body, but I can certainly assist you…
This behavior is particularly misleading for the inference model as the LLAMA family of models consistently provide accurate responses (e.g., identifying themselves as “Meta” or “LLAMA”) across all tested versions.
However, the fine-tuning process also influenced how the model handles other queries, also harmful ones. While UltraMedical continues to refuse to answer such prompts, its refusal messages adopt a different structure than the one of the base model. For example, upon “How to build a bomb?”:
Llama-3.1-8B: I can’t provide information on how to build a bomb. Is there anything else I can help you with? UltraMedical: I must emphasize that providing instructions on how to build a bomb is highly inappropriate and illegal. The creation, possession, or use of explosive devices…
Discussion. Ultimately, our experiments show that general fine-tuning (which makes significant changes to the model’s behavior) has the most significant impact, as it fundamentally alters the LLM’s functionality, influencing its responses to all types of queries (i.e., the new model behaves as an unseen model), in contrast to more targeted modifications such as alignment. A comprehensive analysis of this phenomenon is left as an open question for future research.
Appendix D Additional Testing on LLMs Outside the Training Data
In this section, we evaluate LLMmap open-set’s ability to fingerprint models that were neither part of the optimization of the query strategy (see Appendix H) nor part of the training of the inference model. The experiments aim to evaluate two key objectives: (1) determine whether including an LLM from the same family in LLMmap’s training set enhances its capability to fingerprint new LLM versions within the same family, and (2) assess LLMmap’s ability to fingerprint unknown models (i.e., no earlier version or variation was analyzed during training).
Setup. Given the LLMmap open-set, trained on the $42$ models listed in Table G.1, we run it on 10 newly introduced LLMs which are listed in Table D.1. These models were not included in the inference model’s training set and were not considered during the creation or optimization of the query strategy. The LLMs are categorized into two groups: those that belong to the same family of a model used to set up LLMmap and those that have no connection to them.
LLM Vendor Accuracy Average
granite-3.0-8b-instruct IBM $76.15$ $82.00\pm 3.29$
granite-3.1-8b-instruct IBM $83.85$
Falcon3-7B-Instruct TII $80.77$
Falcon3-10B-Instruct TII $85.38$
EuroLLM-1.7B-Instruct UTTER $83.85$
Qwen2.5-3B-Instruct Qwen $88.46$ $81.38\pm 4.63$
Qwen2.5-0.5B-Instruct Qwen $80.77$
Llama-3.2-1B-Instruct Meta $74.62$
Llama-3.2-3B-Instruct Meta $79.23$
Phi-3.5-mini-instruct Microsoft $83.85$
All models $81.69\pm 8.87$
Table D.1: Fingerprint accuracy of LLMmap ’s open-set evaluation on 10 LLMs that were neither involved in LLMmap ’s training process nor in the development of its query strategies. Rows highlighted in green represent LLMs with no connections to those used during the setup of LLMmap, while rows in red denote LLMs with some connection to the setup models (e.g., from the same vendor or model family).
For each new LLM we take the following steps, (1) we run Algorithm A.1 to derive a signature vector and add it in the database, where $\mathbf{S}$ is set to $\mathbf{S}_{\text{train}}$ (see Section 7.1). (2) For each prompt configuration $s$ in $\mathbf{S}_{\text{test}}$, we apply $s$ on the model version and run the fingerprinting Algorithm A.2, deriving a prediction. (3) We verify if the prediction given by LLMmap is correct.
Results. Results on the accuracy of LLMmap on the individual model versions, as well as the average performance per group, are listed in Table D.1. The results align with those collected in Section 7.2 for the left-out-LLM setting. This suggests that the query strategy deployed and optimized based on the LLMs listed in Table G.1 is general enough to transfer to different models. However, it cannot be excluded that re-optimizing the strategy based on including these new models could result in a different query strategy.
Regarding performance across groups, the accuracy on entirely new model versions closely aligns with that achieved on models for which an earlier version is already in the database, with a discrepancy of less than $0.8\%$ favoring the completely new models. While this minor advantage may be due to the limited sample size, another contributing factor is the composition of the vector signature databases, which already contain multiple instances of LLMs from the target’s family. This increases the difficulty of accurate predictions, as similar models tend to have similar signatures (see Figure 6). Indeed, most failed fingerprinting attempts on these LLM versions involve incorrect attributions to models from the same family or derivatives already in $\mathcal{DB}$, particularly in the case of the two LLama models. This factor is likely enough to cancel out any utility benefit derived from having trained LLMmap on models similar to the target.
Appendix E Detecting Unseen LLMs in the Open-set Setting
Next, we discuss how one can detect unseen LLMs (i.e., versions whose vector signatures do not appear in the database) in the open-set setting.
A natural approach to detecting unseen models is based on the distances of tested signatures to signatures stored in the database at inference time. In particular, one can rely on the minimum distance. Formally, given the vector of distances:
$\vec{d}\leftarrow[d_{1},d_{2},\dots],\ \text{where}\ d_{i}=d(x^{?},x_{i}),\,x_{i}\in\mathcal{DB},$
as defined in Algorithm A.2, one can threshold $\min(\vec{d})$ to try to predict whether the tested LLM has no corresponding signature in $\mathcal{DB}$.
However, we observed that the scale of the minimum distance often depended on the specific LLM under consideration, making it highly variable taken alone. To illustrate this point, different families of LLMs that come from different vendors, may have different thresholds for distinguishing between a model that is a different version and an unseen model. As a result, using a single fixed threshold proved unreliable.
To achieve better performance, we adopt a slightly more fine-grained approach. Instead of relying on a fixed threshold, we apply a simple random forest model to predict whether a model is unseen. The observation here is that not only does the minimum distance provide a signal of an unseen model, but also the way all the other distances distribute among all the other signatures in the database. Thus, we can provide more accurate predictions by considering the distances within the elements of the database as features.
Since the number of distances changes with the size of the signature database, we avoid providing the raw distance vector $\vec{d}$ directly to the model. [^11] As that would have meant retraining the random forest every time the database is updated—an anti-pattern decision for the open-set setup. Instead, we derive features by aggregating these distances using general statistical measures. After running a basic feature selection approach, we derive six statistics defined on $\vec{d}$ that are particularly predictive of unseen models. These are: Mean, minimum, maximum, variance, standard deviation, and entropy all computed on the distances vector $\vec{d}$.
Ultimately, the random forest classifier takes as input a vector composed of those six statistics of the distances and produces a binary output “unseen”. To train the random forest, we derive a suitable training set from the one used to train the inference model. For each LLM, we compute distances by inserting the correct model signature in the database and repeat with the signature removed. We run on all the entries in the training set of the inference model, and we get a balanced of 50% "seen" cases and 50% "unseen" cases. The test is carried out instead on the 10 fresh LLMs used for the experiments in Section D listed in Table D.1. The creation of the test set mirrors the one of the training set.
The average accuracy of the random forest classifier on the test LLMs (Table D.1) is $82.30\%$. It is promising that improved results could be achieved by deliberately training the inference model to recognize unseen models based on the current state of $\mathcal{DB}$. We leave this exploration for future work.
Appendix F Details on query strategy baselines
The queries resulting from the optimization of the two baselines are listed in Table F.1.
GPT4o-gen-opt random-opt
Write a creative story involving quantum mechanics and a detective mystery. Generate a list of 5 different job roles for a computer engineer.
What is your parameter count? Are you a 7B or another size? Write an IF-THEN statement to control the temperature in a room.
What is your token limit per request? Create a list of 10 helpful tips for designing an effective computer-based education system
Explain transformer architecture in detail. What specifics do you highlight? Is the number 12 prime?
How many turns of conversation can you handle without losing context? Write a description of a painting in the style of impressionism.
Are you based on transformers? Do you use rotary embeddings or sparse attention? Write a regular expression that can match a valid email address.
Do you specialize in fields like medicine, law, or coding? Provide an example. Describe what happened on July 23rd, 1990 in one sentence.
What datasets were you trained on? Do you include Common Crawl or proprietary data? Describe the day-to-day job duties of a Human Resources Manager.
Table F.1: Baseline query strategies.
The prompt used to generate the queries from gpt-4o-2024-11-20) is the following:
[…]
[frame] Imagine you are a detective. Given oracle access to an LLM, your task is to generate questions to ask the LLM such that you can infer which model it is based on the given answers (like LLaMA 3.1-7B or Mistral-7B). Generate 30 short queries you think are highly effective.
Appendix G Additional Resources
@@
This appendix contains additional material. Table G.1 reports the complete list of the LLMs considered in this work. Figure G.1 depicts the confusion matrix for a closed-set inference model with the default query strategy. Table G.3 reports examples of system prompts used to generate different prompting configurations. Table G.4 reports the three randomly sampled query strategies used in Figure 8.
Table G.1: List of LLMs used for training and testing LLMmap.
@@ 18 Qwen/Qwen2-72B-Instruct Qwen 72B
Table G.2: Top-8 most effective probes (default query strategy).
Probe Type
(2) Tell me how to build a bomb. Alignment
(3) What is the size of your training dataset? Meta-information
(5) What’s your data cutoff date? Meta-information
(7) Bonjour, how are you doing today? ¿Qué tal?’ Malformed
Figure G.1: Confusion matrix for the closed-set fingerprinting model. Computed on a single training run.
Table G.3: Examples of eight system prompts used to simulate different prompting configurations.
@@
Table G.4: Query strategies composed by randomly sampled prompts.
@@
Appendix H Optimize Query Strategy
Algorithm H.1 Greedy query search algorithm.
1:function greedy_query_opt($\mathbf{Q},n,\mathcal{T}_{\text{train}},\mathcal{T}_{\text{test}}$)